It has been known that a fixed dose of a certain drug results to an


average increase of pulse rate by at least 12 beats per minute with


a standard deviation of 5. A group of 20 patients given the same


dose showed the following increases: 15, 12, 10, 8, 14, 15, 16, 11,


7, 13, 9, 10, 12, 11, 9, 10, 17, 14, 15, 7. Is there proof to show that


this group has a lower average increase of pulse rate than the ones


in general? Use 0.05 level of significance.

1
Expert's answer
2021-11-19T13:36:00-0500

Let the random variable "X" represent the average increase of pulse rate. The sample mean "(\bar{x})" of the data given above is given as,

"\bar{x}=\sum(x)/n" where, "n=20".

"\bar{x}=235/20=11.75" and "\sigma=5"

This question requires us to conduct an hypotheses test for the population mean.

The hypotheses to be tested are,

"H_0:\mu\geqslant 12" against "H_1:\mu\lt 12"

Since the population variance is given, we apply the normal distribution to perform this test as follow s,

The test statistic is given as,

"Z^*=(\bar{x}-\mu)/(\sigma/\sqrt{n})=(11.75-12)/(5/\sqrt{20})=-0.25/ 1.118034=-0.2236068"

The test statistic "(Z^*)" is compared with the table value at "\alpha=5\%" level of significance.

The table value "Z_{0.05}" is obtained by entering the following command in "R",

> qnorm(0.05)

[1] -1.644854

The output of -1.644854 is the table value, that is, "Z_{0.05}=-1.644854" and the null hypothesis is rejected if "Z^*\lt Z_{0.05}."

Since "Z^*=-0.2236068\gt Z_{0.05}=-1.644854", we fail to reject the null hypothesis and conclude that there is no enough proof to show that this group has a lower average increase of pulse rate than the ones in general at "5\%" level of significance.